# spacing=nonfrench

### v1

@@ Abstract.
Fast and scalable metadata management across multiple metadata servers is crucial for distributed file systems to handle numerous files and directories. Client-side caching of frequently accessed metadata can mitigate server loads, but incurs significant overhead and complexity in maintaining cache consistency when the number of clients increases. We propose FMCache, an in-switch file-system metadata caching framework that leverages programmable switches to serve file-system metadata requests from multiple clients directly in the switch data plane. Unlike prior in-switch key-value caching approaches, FMCache addresses file-system-specific path dependencies under stringent switch resource constraints. We implement FMCache atop Hadoop HDFS and evaluate it on a Tofino-switch testbed using real-world file-system metadata workloads. FMCache achieves up to 181.6% higher throughput than vanilla HDFS and complements client-side caching with additional throughput gains of up to 139.6%. It also incurs low latencies and limited switch resource usage.
1. Introduction
@@
Scaling file-system metadata management across multiple metadata servers is crucial for distributed file systems to handle billions of files and directories. Field studies, both classical (Roselli et al., 2000; Harter et al., 2012) and recent (Wang et al., 2023), indicate that metadata operations dominate file-system requests. For example, 67-96% of file-system requests in Baidu AI Cloud are metadata-related (Wang et al., 2023). As the number of files and directories grows, especially in workloads dominated by small files (Carns et al., 2009; Harter et al., 2014), metadata operations, such as inode lookups, permission checks, and directory traversals, become performance bottlenecks.
@@
Caching frequently accessed metadata on the client side is a common strategy to mitigate loads on metadata servers. File-system metadata access patterns in practice tend to be skewed, with a small fraction of files and directories being accessed far more frequently than others (Leung et al., 2008; Abad et al., 2012; Ananthanarayanan et al., 2011); for example, less than 3% of files account for 34-39% of requests in Yahoo’s HDFS clusters (Abad et al., 2012). However, client-side caching incurs significant overhead in maintaining cache consistency across numerous clients, as metadata updates require notifications to all clients to invalidate cached metadata. Existing client-side caching approaches (Welch et al., 2008; Ren et al., 2014; Li et al., 2017a; Lv et al., 2022) cache limited metadata information to mitigate path resolution overhead (§2.1) but rely on metadata servers for accessing metadata contents, so the performance gains are limited.
@@
Programmable switches (Bosshart et al., 2013) offer a promising alternative by enabling in-switch caching across multiple clients and leveraging a centralized view of requests, so as to eliminate switch-to-server transmissions and server-side processing. While prior work has explored in-switch caching for key-value stores (Li et al., 2016; Jin et al., 2017; Liu et al., 2019; Sheng et al., 2023; Li et al., 2020; Eldakiky et al., 2021; Lee et al., 2021), in-switch caching for file-system metadata poses unique challenges not directly addressed before. First, file-system pathnames, unlike fixed-length keys, have large and variable sizes and cannot readily fit into limited switch resources. Second, accessing a file’s metadata requires accessing its internal directories’ metadata, thereby exacerbating switch resource demands for caching multiple levels of metadata. Third, cache lookups for the metadata of files and directories require multiple iterations under the strict switch programming model, thereby complicating concurrent cache updates and lookups while maintaining consistency and performance.
@@
We propose FMCache, an in-switch file-system metadata caching framework tailored for distributed file systems under skewed, read-intensive workloads. FMCache extends in-switch key-value caches by specifically addressing file-system path dependencies through techniques with performance and correctness guarantees: (i) path-aware cache management to account for path dependencies in cache admission and eviction; (ii) multi-level read-write locking to enable high-performance concurrent cache lookups and updates; and (iii) local hash collision resolution to efficiently and correctly map variable-length paths into fixed-length keys.
@@
We implemented FMCache in P4 (Bosshart et al., 2014) and compiled it into the Tofino switch chipset (tof, [n. d.]; swi, [n. d.]). We integrated FMCache with Hadoop HDFS (hdf, [n. d.]a), while preserving HDFS semantics. We also implemented a state-of-the-art client-side caching approach (Ren et al., 2014; Lv et al., 2022). Our evaluation on four real-world workloads (Ren et al., 2014; Lv et al., 2022; Xu et al., 2024) shows that under 128 simulated metadata servers, FMCache achieves up to 181.6% higher throughput than vanilla HDFS, while client-side caching coupled with FMCache achieves up to 139.6% higher throughput than without FMCache.
2. Background and Motivation
2.1. File-System Metadata Management
@@
In this work, we focus on Hadoop Distributed File System (HDFS) (Shvachko et al., 2010), where a namenode manages the namespace and metadata. HDFS supports various metadata operations for files (e.g., create, delete, open, and close), directories (e.g., mkdir and rmdir), attributes (e.g., chmod, chown, and utime), and data organization (i.e., rename, readdir, and stat). Metadata operations rely on path resolution, which parses and traverses each level of a path to verify metadata for existence and permissions. For scalability, HDFS employs Router-Based Federation (RBF) (hdf, [n. d.]b) to distribute namespace and metadata management across multiple namenodes.
Figure 1. Data plane of a programmable switch.
2.2. Programmable Switches
Programmable switches (Bosshart et al., 2013) enable customized packet processing tasks beyond traditional packet forwarding. We focus on Tofino switches (tof, [n. d.]), which are programmed using the P4 language (Bosshart et al., 2014) and built on the Protocol-Independent Switch Architecture (PISA) based on the Reconfigurable Match-Action Table (RMT) paradigm (Bosshart et al., 2013). We target the generic RMT architecture without relying on any Tofino-specific features, so our design is extensible to other RMT-based switch platforms (bro, [n. d.]; Chole et al., 2017; hua, [n. d.]). A programmable switch comprises a data plane and a control plane. The data plane, as shown in Figure 1, performs packet forwarding via multiple ingress and egress pipelines. Packets enter an ingress pipeline and are forwarded to an egress pipeline via a traffic manager. The control plane manages the data plane by specifying packet processing rules.
Each ingress or egress pipeline processes packets via a series of stages, each of which contains match-action tables (MATs) that execute processing logic. A parser converts packet headers into packet header vectors (PHVs), processed by MATs with on-chip ALUs across stages. Each stage can only access limited SRAM, typically with tens of memory blocks. The switch programming model imposes strict constraints: each memory block can be accessed at most once per PHV traversal, and stages cannot access memory blocks of other stages. After processing, a deparser reconstructs packets. To process packets in multiple iterations, switches support recirculation, which redirects packets from an egress pipeline back to an ingress pipeline. Recirculation should be cautiously used, as it consumes extra switch resources and slows down packet forwarding. In this work, we leverage recirculation for read requests that require in-switch path resolution (§4.1) and write requests that await locks (§5.2).
2.3. Challenges
@@
Challenge 1: Constrained switch resources. Switches have limited resources and complicate caching implementation. For example, Tofino switches (tof, [n. d.]) provide only 12 stages per ingress or egress pipeline, each processing up to 16 bytes due to ALU word size restrictions. The PHV size is capped (e.g., 768 bytes in Tofino switches (tof, [n. d.])), and all pipelines share limited SRAM. For comparisons, NetCache (Jin et al., 2017) allocates SRAM for access frequency tracking and supports only a maximum value size of 128 bytes for key-value records.
File-system metadata caching is particularly resource-intensive. HDFS file or directory names can reach 255 bytes (def, [n. d.]) and aggravate SRAM and PHV overhead. Path resolution for a path (e.g., /a/b.txt) requires accessing metadata for its ancestors (i.e., / and /a) (§2.1). Caching metadata for all paths and their ancestors is desirable but challenging given limited SRAM. Splitting path processing across multiple levels requires careful synchronization across stages. To mitigate switch resource usage, existing in-switch key-value caches (Jin et al., 2017; Liu et al., 2019; Sheng et al., 2023) offload cache admission and eviction to a centralized controller, but excessive switch-to-controller communications incur high latencies.
@@
3. Design Overview
3.1. Goals and Assumptions
@@
FMCache is an in-switch metadata caching framework for distributed file systems, aiming to achieve high throughput and low latencies for file-system metadata requests across multiple metadata servers. It is currently designed for HDFS (§2.1), supporting all HDFS metadata operations via its interface (§7) and preserving HDFS semantics. It is also extensible to other distributed file systems via their respective APIs.
@@
FMCache targets read-intensive file-system metadata workloads, as observed in prior studies (Ren et al., 2014; Lv et al., 2022; Xu et al., 2024). For example, in a LinkedIn HDFS cluster, 84% of 145 million metadata operations are lookups, with only 9% creates and 7% updates (Ren et al., 2014); in Alibaba’s Pangu file system, over 60% of metadata operations are reads (Lv et al., 2022). To ensure consistency, FMCache adopts write-through caching (as in prior in-switch key-value caches (Li et al., 2016; Liu et al., 2017; Jin et al., 2017; Liu et al., 2019)) to update both the in-switch cache and server-side metadata in writes before acknowledging clients. Also, since path resolution always starts at the root, FMCache assumes the root directory’s metadata is persistently cached, with its permissions unchanged and the root not being deleted (Niazi et al., 2017).
3.2. Architecture and Design Roadmap
Figure 2. FMCache’s architecture.
@@ 3.2. Architecture and Design Roadmap
Figure 2 depicts FMCache’s architecture, which comprises both control-plane and data-plane components. In the control plane, a controller manages in-switch cache admission and eviction, while in the data plane, multiple clients send (receive) file-system metadata requests (responses) via the switch to (from) multiple metadata servers (or servers for short). FMCache employs three techniques to enable efficient in-switch file-system metadata caching.
@@
Path-aware cache management (§4). FMCache caches file-system metadata in the switch data plane. FMCache hashes variable-length paths into fixed-length keys and manages file-system metadata in key-value pairs as in existing in-switch key-value caches (Li et al., 2016; Jin et al., 2017; Liu et al., 2019; Sheng et al., 2023). But unlike in-switch key-value caches, FMCache addresses path dependencies across hierarchical levels during cache admission and eviction.
@@
Multi-level read-write locking (§5). To synchronize concurrent metadata operations, FMCache adopts multi-level read-write locking using (i) lock counters as read-write locks for different path levels and (ii) validation flags for verifying the validity of cached paths. It ensures reliable locking and unlocking during path resolution even in unreliable networks.
@@
Local hash collision resolution (§6). Mapping paths to fixed-size keys in the switch data plane can cause hash collisions, leading to incorrect metadata retrieval. For example, if two paths /a and /b share the same hash and /a is cached, a request for /b can erroneously return /a’s metadata. FMCache proposes a local hash collision resolution approach, where the controller assigns unique values, called tokens, to hash-colliding paths and distributes the tokens to clients, servers, and the switch, so as to allow local resolution of hash collisions without the controller’s intervention for every request. FMCache ensures no incorrect metadata retrieval.
4. Path-Aware Cache Management
@@
FMCache incorporates path awareness into cache admission and eviction by caching frequently accessed paths and their ancestors, so as to mitigate cache misses during path resolution. This ensures that if a path is cached, its ancestors are also cached, thereby allowing the reuse of cached metadata across different paths with the common ancestors.
4.1. Path Representation
@@
FMCache maps a file-system pathname into fixed-size hash keys for efficient switch operations. For a read operation, a FMCache client partitions a path into multiple levels (including the root) and computes a hash for each level. For example, /a/b/c.txt is partitioned into /, /a, /a/b, and /a/b/c.txt, with each level being hashed. To avoid redundant computations, the hash of the root directory / is pre-computed and cached on the client side. For a write operation, the client computes a hash for the complete path without partitioning, as all writes are forwarded to the server (that holds the namespace) for path resolution under write-through caching (§3.1). Currently, FMCache uses the first 64 bits of a 128-bit MD5 hash as the hash value.
@@
FMCache caches file-system metadata as key-value records in the switch. Each record is identified by its hash key, and its value contains file or directory metadata. Files and directories have common metadata fields (i.e., type, permissions, owner/group IDs, and timestamps); while files additionally include the size and replication factor fields. Each file’s metadata has 40 bytes, and each directory’s has 24 bytes.
@@
FMCache manages cache across multiple stages in the switch. It leverages MATs across two stages in a single ingress pipeline to perform cache lookups based on hash keys, so as to avoid cross-pipeline synchronization (§5) and enforce logical dependency (i.e., cache hit/miss status is resolved before metadata retrieval can proceed). Since packets are strictly processed in ingress-to-egress order, this placement enables fast cache status checks. The switch manages metadata in 32 registers across eight stages in the egress pipelines (§7). To simplify routing, the switch caches metadata in the egress pipeline that is connected to the corresponding server.
@@
Clients issue metadata requests, including hash keys and full pathnames, to the switch, which reports frequently accessed pathnames to the controller during cache admission (§4.2). For reads, the switch performs path resolution by issuing cache lookups and permission checks for each level, starting from the root, and recirculates the request to the ingress pipeline for each next level. For a cache hit (i.e., the metadata for the path and all its ancestors are cached), the switch returns the metadata if all levels’ permission checks pass, or an error if any check fails. For a cache miss, the switch forwards the request to the server. Due to limited stages and memory access constraints (§2.2), FMCache cannot perform path resolution for multiple path levels in a single pass, but instead leverages recirculation for path resolution while limiting overhead (§8.2).
4.2. Cache Admission and Eviction Workflows

### v2

# Fletch: File-System Metadata Caching in Programmable Switches

@@ Abstract
Fast and scalable metadata management across multiple metadata servers is crucial for distributed file systems to handle numerous files and directories. Client-side caching of frequently accessed metadata can mitigate server loads, but incurs significant overhead and complexity in maintaining cache consistency when the number of clients increases. We explore caching in programmable switches by serving file-system metadata requests from multiple clients on the switch data plane. Despite prior efforts on in-switch key-value caching, they fail to address the path dependencies specific to file-system semantics. We propose Fletch, an in-switch file-system metadata caching framework that leverages programmable switches to serve file-system metadata requests from multiple clients directly in the switch data plane. Unlike prior in-switch key-value caching approaches, Fletch addresses file-system-specific path dependencies under stringent switch resource constraints. We implement Fletch atop Hadoop HDFS and evaluate it on a Tofino-switch testbed using real-world file-system metadata workloads. Fletch achieves up to 181.6% higher throughput than vanilla HDFS and complements client-side caching with additional throughput gains of up to 139.6%. It also incurs low latencies and limited switch resource usage.
I Introduction
@@
Scaling file-system metadata management across multiple metadata servers is crucial for distributed file systems to handle billions of files and directories. Field studies, both classical [roselli00, harter12] and recent [wang23], indicate that metadata operations dominate file-system requests. For example, 67-96% of file-system requests in Baidu AI Cloud are metadata-related [wang23]. As the number of files and directories grows, especially in workloads dominated by small files [carns09, harter14], metadata operations, such as inode lookups, permission checks, and directory traversals, become performance bottlenecks.
@@
Caching frequently accessed metadata on the client side is a common strategy to mitigate loads on metadata servers. File-system metadata access patterns in practice tend to be skewed, with a small fraction of files and directories being accessed far more frequently than others [leung08, abad12, ananthanarayanan11]; for example, less than 3% of files account for 34-39% of requests in Yahoo’s HDFS clusters [abad12]. However, client-side caching incurs significant overhead in maintaining cache consistency across numerous clients, as metadata updates require notifications to all clients to invalidate cached metadata. Existing client-side caching approaches [welch08, ren14, li17, lv22] cache limited metadata information to mitigate path resolution overhead (§II-A) but rely on metadata servers for accessing metadata contents, so the performance gains are limited.
@@
Programmable switches [bosshart13] offer a promising alternative by enabling in-switch caching across multiple clients and leveraging a centralized view of requests, so as to eliminate switch-to-server transmissions and server-side processing. While prior work has explored in-switch caching for key-value stores [li16, jin17, liu19, sheng25, li20, eldakiky20, lee21], in-switch caching for file-system metadata poses unique challenges not directly addressed before. First, file-system pathnames, unlike fixed-length keys, have large and variable sizes and cannot readily fit into limited switch resources. Second, accessing a file’s metadata requires accessing its internal directories’ metadata, thereby exacerbating switch resource demands for caching multiple levels of metadata. Third, cache lookups for the metadata of files and directories require multiple iterations under the strict switch programming model, thereby complicating concurrent cache updates and lookups while maintaining consistency and performance.
@@
We propose Fletch, an in-switch file-system metadata caching framework tailored for distributed file systems under skewed, read-intensive workloads. Fletch extends in-switch key-value caches by specifically addressing file-system path dependencies through techniques with performance and correctness guarantees: (i) path-aware cache management to account for path dependencies in cache admission and eviction; (ii) multi-level read-write locking to enable high-performance concurrent cache lookups and updates; and (iii) local hash collision resolution to efficiently and correctly map variable-length paths into fixed-length keys.
@@
We implemented Fletch in P4 [bosshart14] and compiled it into the Tofino switch chipset [tofino, switchthpt]. We integrated Fletch with Hadoop HDFS [hdfs], while preserving HDFS semantics. We also implemented a state-of-the-art client-side caching approach [ren14, lv22]. Our evaluation on four real-world workloads [ren14, lv22, xu24] shows that under 128 simulated metadata servers, Fletch achieves up to 181.6% higher throughput than vanilla HDFS, while client-side caching coupled with Fletch achieves up to 139.6% higher throughput than without Fletch. We release the source code of Fletch at https://github.com/calanquee/Fletch.
II Background and Motivation
II-A File-System Metadata Management
@@
In this work, we focus on Hadoop Distributed File System (HDFS) [shvachko10], where a namenode (i.e., metadata server) manages the namespace and metadata. HDFS supports various metadata operations for files (e.g., create, delete, open, and close), directories (e.g., mkdir and rmdir), attributes (e.g., chmod, chown, and utime), and data organization (i.e., rename, readdir, and stat). Metadata operations rely on path resolution, which parses and traverses each level of a path to verify metadata for existence and permissions. For scalability, HDFS employs Router-Based Federation (RBF) [hdfsrbf] to distribute namespace and metadata management across multiple namenodes.
Figure 1: Data plane of a programmable switch.
II-B Programmable Switches
Programmable switches [bosshart13] enable customized packet processing tasks beyond traditional packet forwarding. We focus on Tofino switches [tofino], which are programmed using the P4 language [bosshart14] and built on the Protocol-Independent Switch Architecture (PISA) based on the Reconfigurable Match-Action Table (RMT) paradigm [bosshart13]. We target the generic RMT architecture without relying on any Tofino-specific features, so our design is extensible to other RMT-based switch platforms [broadcomswitch, chole17, huaweiswitch]. A programmable switch comprises a data plane and a control plane. The data plane, as shown in Figure 1, performs packet forwarding via multiple ingress and egress pipelines. Packets enter an ingress pipeline and are forwarded to an egress pipeline via a traffic manager. The control plane manages the data plane by specifying packet processing rules.
Each ingress or egress pipeline processes packets via a series of stages, each of which contains match-action tables (MATs) that execute processing logic. A parser converts packet headers into packet header vectors (PHVs), processed by MATs with on-chip ALUs across stages. Each stage can only access limited SRAM, typically with tens of memory blocks. The switch programming model imposes strict constraints: each memory block can be accessed at most once per PHV traversal, and stages cannot access memory blocks of other stages. After processing, a deparser reconstructs packets. To process packets in multiple iterations, switches support recirculation, which redirects packets from an egress pipeline back to an ingress pipeline. Recirculation should be cautiously used, as it consumes extra switch resources and slows down packet forwarding. In this work, we leverage recirculation for read requests that require in-switch path resolution (§IV-A) and write requests that await locks (§V-B).
II-C Challenges
@@
Challenge 1: Constrained switch resources. Switches have limited resources and complicate caching implementation. For example, Tofino switches [tofino] provide only 12 stages per ingress or egress pipeline, each processing up to 16 bytes due to ALU word size restrictions. The PHV size is capped (e.g., 768 bytes in Tofino switches [tofino]), and all pipelines share limited SRAM. For comparison, NetCache [jin17] allocates SRAM for access frequency tracking and supports only a maximum value size of 128 bytes for key-value records.
File-system metadata caching is particularly resource-intensive. HDFS file or directory names can reach 255 bytes [defaultsetting] and aggravate SRAM and PHV overhead. Path resolution for a path (e.g., /a/b.txt) requires accessing metadata for its ancestors (i.e., / and /a) (§II-A). Caching metadata for all paths and their ancestors is desirable but challenging given limited SRAM. Splitting path processing across multiple levels requires careful synchronization across stages. To mitigate switch resource usage, existing in-switch key-value caches [jin17, liu19, sheng25] offload cache admission and eviction to a centralized controller, but excessive switch-to-controller communications incur high latencies.
@@
III Design Overview
III-A Goals and Assumptions
@@
Fletch is an in-switch metadata caching framework for distributed file systems, aiming to achieve high throughput and low latencies for file-system metadata requests across multiple metadata servers. It is currently designed for HDFS (§II-A), supporting all HDFS metadata operations via its interface (§VII) and preserving HDFS semantics. It is also extensible to other distributed file systems via their respective APIs.
@@
Fletch targets read-intensive file-system metadata workloads, as observed in prior studies [ren14, lv22, xu24]. For example, in a LinkedIn HDFS cluster, 84% of 145 million metadata operations are lookups, with only 9% creates and 7% updates [ren14]; in Alibaba’s Pangu file system, over 60% of metadata operations are reads [lv22]. To ensure consistency, Fletch adopts write-through caching (as in prior in-switch key-value caches [li16, liu17, jin17, liu19]) to update both the in-switch cache and server-side metadata in writes before acknowledging clients. With write-through caching, Fletch avoids possible data loss caused by switch failures as metadata servers always possess the latest file-system metadata. Also, since path resolution always starts at the root, Fletch assumes the root directory’s metadata is persistently cached, with its permissions unchanged and the root not being deleted [niazi17].
III-B Architecture and Design Roadmap
Figure 2: Fletch’s architecture.
@@ III-B Architecture and Design Roadmap
Figure 2 depicts Fletch’s architecture, which comprises both control-plane and data-plane components. In the control plane, a controller manages in-switch cache admission and eviction, while in the data plane, multiple clients send (receive) file-system metadata requests (responses) via the switch to (from) multiple metadata servers (or servers for short). Fletch employs three techniques to enable efficient in-switch file-system metadata caching.
@@
Path-aware cache management (§IV). Fletch caches file-system metadata in the switch data plane. Fletch hashes variable-length paths into fixed-length keys and manages file-system metadata in key-value pairs as in existing in-switch key-value caches [li16, jin17, liu19, sheng25]. But unlike in-switch key-value caches, Fletch addresses path dependencies across hierarchical levels during cache admission and eviction.
@@
Multi-level read-write locking (§V). To synchronize concurrent metadata operations, Fletch adopts multi-level read-write locking using (i) lock counters as read-write locks for different path levels and (ii) validation flags for verifying the validity of cached paths. It ensures reliable locking and unlocking during path resolution even in unreliable networks.
@@
Local hash collision resolution (§VI). Mapping paths to fixed-size keys in the switch data plane can cause hash collisions, leading to incorrect metadata retrieval. For example, if two paths /a and /b share the same hash and /a is cached, a request for /b can erroneously return /a’s metadata. Fletch proposes a local hash collision resolution approach, where the controller assigns unique values, called tokens, to hash-colliding paths and distributes the tokens to clients, servers, and the switch, so as to allow local resolution of hash collisions without the controller’s intervention for every request. Fletch ensures no incorrect metadata retrieval.
IV Path-Aware Cache Management
@@
Fletch incorporates path awareness into cache admission and eviction by caching frequently accessed paths and their ancestors, so as to mitigate cache misses during path resolution. This ensures that if a path is cached, its ancestors are also cached, thereby allowing the reuse of cached metadata across different paths with the common ancestors.
IV-A Path Representation
@@
Fletch maps a file-system pathname into fixed-size hash keys for efficient switch operations. For a read operation, a Fletch client partitions a path into multiple levels (including the root) and computes a hash for each level. For example, /a/b/c.txt is partitioned into /, /a, /a/b, and /a/b/c.txt, with each level being hashed. To avoid redundant computations, the hash of the root directory / is pre-computed and cached on the client side. For a write operation, the client computes a hash for the complete path without partitioning, as all writes are forwarded to the server (that holds the namespace) for path resolution under write-through caching (§III-A). Currently, Fletch uses the first 64 bits of a 128-bit MD5 hash as the hash value; note that MD5 is used for fast non-cryptographic hashing; collision resistance is addressed separately via tokens (§VI).
@@
Fletch caches file-system metadata as key-value records in the switch. Each record is identified by its hash key, and its value contains file or directory metadata. Files and directories have common metadata fields (i.e., type, permissions, owner/group IDs, and timestamps); while files additionally include the size and replication factor fields. Each file’s metadata has 40 bytes, and each directory’s has 24 bytes.
@@
Fletch manages cache across multiple stages in the switch. It leverages MATs across two stages in a single ingress pipeline to perform cache lookups based on hash keys, so as to avoid cross-pipeline synchronization (§V) and enforce logical dependency (i.e., cache hit/miss status is resolved before metadata retrieval can proceed). Since packets are strictly processed in ingress-to-egress order, this placement enables fast cache status checks. The switch manages metadata in 32 registers across eight stages in the egress pipelines (§VII). To simplify routing, the switch caches metadata in the egress pipeline that is connected to the corresponding server.
@@
Clients issue metadata requests, including hash keys and full pathnames, to the switch, which reports frequently accessed pathnames to the controller during cache admission (§IV-B). For reads, the switch performs path resolution by issuing cache lookups and permission checks for each level, starting from the root, and recirculates the request to the ingress pipeline for each next level. For a cache hit (i.e., the metadata for the path and all its ancestors are cached), the switch returns the metadata if all levels’ permission checks pass, or an error if any check fails. For a cache miss, the switch forwards the request to the server. Due to limited stages and memory access constraints (§II-B), Fletch cannot perform path resolution for multiple path levels in a single pass, but instead leverages recirculation for path resolution while limiting overhead (§VIII-B).
IV-B Cache Admission and Eviction Workflows